1. Yoplait has new line of Greek yogurt in a 9 oz size. The Marketing Managers plan to test market the product in one small market for three weeks. They has calculated the cost and . projected sales for each of the different flavors: Plain cost is .55 each with a projected sales of 10,000 units, Strawberry cost per unit is .60 with a projected sales of 13,000 units, Blueberry cost are .68 per unit with a projected sales of 9,000 units. Each yogurt product will retail at the same price. Using the average cost pricing and a 35% markup what will Yoplait price these product at the retail level ______________
Answer #1

First find out the total cost :

Total cost = total cost of plain + total cost of strawberry + total cost of blueberry

Each flavors total cost formula= cost per unit × projected sales

Total cost of plain= 0.55 × 10,000= 5500

Total cost of strawberry = 0.6× 13000= 7800

Total cost of blueberry = 0.68× 9000= 6120

Total cost = 5500+ 7800+ 6120= 19420

Average cost = total cost / total units

Total units = plain units + strawberry units + blueberry units

Total units = 10000+ 13000+9000= 32000

Average cost= 19420/32000= 0.61

Selling price = average cost + markup

Markup = 35% × average cost

Selling price = 0.61 + (0.35× 0.61)

Selling price = 0.61+0.21= 0.82

0.82 is a yoplait price these products at the retail level.